OpenStudy (anonymous):

How to do expression like: 3 square root of 2?

OpenStudy (anonymous):

\[\large \large3\sqrt{2} = ?\]

OpenStudy (jdoe0001):

hmmm to do ? what do you mean?

OpenStudy (anonymous):

Like how to find the value of 3 square root of 2.

OpenStudy (anonymous):

Sorry that I wasn't clear enough.

OpenStudy (anonymous):

The value will only represent an approximation

OpenStudy (jdoe0001):

well... using a calculator I gather, \(\bf \sqrt{2}\approx 1.414213562\)

OpenStudy (anonymous):

I've seen people finding the value of \[\large \large \frac{ 1 }{ 2 } (3\sqrt{2})(3\sqrt{2})\]

OpenStudy (jdoe0001):

yes, because you have a pair of those... so that changes things some

OpenStudy (anonymous):

Alright, how do you do that expression?

OpenStudy (anonymous):

How do you find the value of that expression*

OpenStudy (jdoe0001):

\(\bf \cfrac{1}{2}(3\sqrt{2})(3\sqrt{2})\\ \quad \\ \quad \\ \color{blue}{(3\sqrt{2})(3\sqrt{2})\implies (3\sqrt{2})^2\implies 3^2(\sqrt{2})^2\implies 9\sqrt{2^2}\implies 9\times 2}\\ \quad \\ thus\\ \quad \\ \cfrac{1}{2}(3\sqrt{2})(3\sqrt{2})\implies \cfrac{1}{2}\times 9\times 2\)

OpenStudy (anonymous):

Oh. Thank you! Btw, how do you do that drawing?

OpenStudy (jdoe0001):

drawing?

OpenStudy (jdoe0001):

\(\bf \cfrac{1}{2}\times 9\times 2\implies \cfrac{9\times \cancel{2}}{\cancel{2}}\implies 9\)

OpenStudy (anonymous):

How do you show me that steps? Did you use the equation button?

OpenStudy (jdoe0001):

yes and no, I us the same .... code to render it, Latex, I do not use the editor here, I use my own editor, why? is far faster and flexible =)

OpenStudy (anonymous):

What's latex? o-o

OpenStudy (anonymous):

So you basically typed out the code?

OpenStudy (jdoe0001):

a language mainly for writing equations or math glyphs or symbols

OpenStudy (jdoe0001):

well, kinda.. yes, I use my editor to do a good chunk of it, I could have typed it, yes, but using my editor is much faster =)

OpenStudy (anonymous):

How did you get the editor?

OpenStudy (jdoe0001):

well... you can just search around for a LaTex editor, there are some online, many free

OpenStudy (anonymous):

Alright, I'll do that. Thanks. :)

OpenStudy (jdoe0001):

yw